# On discount functions for economic model predictive control without terminal conditions

### v1

@@ Appendix A Technical Lemmas
$\bigg{|}\sum_{j\in\mathcal{J}}\sum_{k\in\mathcal{I}_{j}}\left({\textstyle\beta\big{(}\frac{k-1}{N}\big{)}}-{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\right)g\left(x_{x}^{\star}(k),u_{x}^{\star}(k)\right)\bigg{|}\leq(\sqrt[4]{N^{3}}+1)p^{\star}\bar{g}\frac{L_{\beta}}{N}=:\delta_{4}(N)$
@@
where continuity of $g$ and compactness of $\mathbb{X}\times\mathbb{U}$ imply $\bar{g}=\max_{(x,u)\in\mathbb{X}\times\mathbb{U}}|g(x,u)|$ exists and is finite. Note that $\delta_{4}\in\mathcal{L}$. Further, with the same reasons there exists $\overline{\alpha}_{g}\in\mathcal{K}_{\infty}$ that satisfies for all $k\in\mathbb{I}_{[0,p^{\star}-1]}$ and all $(x,u)\in\mathbb{X}\times\mathbb{U}$
$|g(x,u)-g({\Pi^{\star}}(k))|$ $\leq\overline{\alpha}_{g}(\|(x,u)-{\Pi^{\star}}(k)\|).$
Next, define the complement of $\mathcal{J}$ by $\bar{\mathcal{J}}:=\mathbb{I}_{[0,\lfloor(N-1)/p^{\star}\rfloor]}\setminus\mathcal{J}$, and compute
@@
$\Sigma_{1}$ $\leq\bigg{|}g^{\star}-\sum_{j\in\bar{\mathcal{J}}}\sum_{k\in\mathcal{I}_{j}}\left({\textstyle\beta\big{(}\frac{k-1}{N}\big{)}}-{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\right)g\left(x_{x}^{\star}(k),u_{x}^{\star}(k)\right)\bigg{|}+\delta_{4}(N)$
@@
$\leq\bigg{|}g^{\star}-\sum_{j\in\bar{\mathcal{J}}}\sum_{k\in\mathcal{I}_{j}}\left({\textstyle\beta\big{(}\frac{k-1}{N}\big{)}}-{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\right)g\left({\Pi^{\star}}(i_{k})\right)\bigg{|}+\underbrace{L_{\beta}\overline{\alpha}_{g}(\sigma(N))+\delta_{4}(N)}_{=:\delta_{5}(N)}$
@@
where $i_{k}\in\mathbb{I}_{[0,p^{\star}-1]}$ is such that $\|{\Pi^{\star}}(i_{k})-(x_{x}^{\star}(k),u_{x}^{\star}(k))\|\leq\sigma(N)$ and $\delta_{5}\in\mathcal{L}$. First, we consider $N\geq N_{0}$ and choose $N_{0}\in\mathbb{N}$ large enough such that $\sigma(N_{0})\leq\bar{\varepsilon}$ with $\bar{\varepsilon}$ from [13, Lemma 15]. Then this Lemma guarantees that the points in $\mathcal{I}_{j}$ follow the orbit, since they are inside the $\sigma(N)$-neighborhood of ${\Pi^{\star}}$, i.e., $i_{k+1}=[i_{k}+1]_{p^{\star}}$ for all $k\in\mathcal{I}_{j}$. Hence, the $p^{\star}$ elements of $\{i_{k}|k\in\mathcal{I}_{j}\}$ are exactly $[0,p^{\star}-1]$ and for each $p\in[0,p^{\star}-1]$ and $j\in\bar{\mathcal{J}}$ there exists a $k_{p,j}\in\mathcal{I}_{j}$ such that $i_{k_{p,j}}=p$. Therefore we can continue as follows
@@
$\Sigma_{1}$ $\leq\bigg{|}\sum_{p=0}^{p^{\star}-1}g({\Pi^{\star}}(p))\bigg{(}\frac{1}{p^{\star}}-\sum_{j\in\bar{\mathcal{J}}}\left({\textstyle\beta\big{(}\frac{k_{p,j}-1}{N}\big{)}}-{\textstyle\beta\big{(}\frac{k_{p,j}}{N}\big{)}}\right)\bigg{)}\bigg{|}+\delta_{5}(N)$
@@
$\leq\sum_{p=0}^{p^{\star}-1}\left|g({\Pi^{\star}}(p))\right|\bigg{|}\frac{1}{p^{\star}}-\sum_{j\in\bar{\mathcal{J}}}\left({\textstyle\beta\big{(}\frac{k_{p,j}-1}{N}\big{)}}-{\textstyle\beta\big{(}\frac{k_{p,j}}{N}\big{)}}\right)\bigg{|}+\delta_{5}(N).$ (14)
@@
$\sum_{j\in\bar{\mathcal{J}}}\left({\textstyle\beta\big{(}\frac{k_{p,j}-1}{N}\big{)}}-{\textstyle\beta\big{(}\frac{k_{p,j}}{N}\big{)}}\right)\leq\sum_{j\in\bar{\mathcal{J}}}-\frac{1}{N}\inf_{\xi\in\left[\frac{k_{p,j}-1}{N},\frac{k_{p,j}}{N}\right]}\beta^{\prime}(\xi)$
@@
$\frac{1}{p^{\star}}-\delta_{6}(N)\leq\sum_{j\in\bar{\mathcal{J}}}\left({\textstyle\beta\big{(}\frac{k_{p,j}-1}{N}\big{)}}-{\textstyle\beta\big{(}\frac{k_{p,j}}{N}\big{)}}\right)\leq\frac{1}{p^{\star}}+\delta_{6}(N).$
@@
Using this in (14) and defining $\delta_{1}(N):=\delta_{5}(N)+\sum_{p=0}^{p^{\star}-1}\left|g({\Pi^{\star}}(p))\right|(\delta_{6}(N))$, which satisfies $\delta_{1}\in\mathcal{L}$, we obtain the desired result for $N\geq N_{0}$. For $N<N_{0}$ we know that $\Sigma_{1}\leq|g^{\star}|+\bar{g}$ and thus by redefining $\delta_{1}(N):=\max\{\delta_{1}(N_{0}),|g^{\star}|+\bar{g}\}$ for $N<N_{0}$ the result follows for all $N\in\mathbb{N}$ as still $\delta_{1}\in\mathcal{L}$.
Lemma A.5 (Almost optimal candidate solution).
$V_{N}^{\beta}(x_{x}^{\star}(1))-V_{N}^{\beta}(x)\leq-\ell(x,u_{x}^{\star}(0))+\ell^{\star}+\delta_{2}(N).$ (15)
Proof A.6.
@@
To begin with, we construct a candidate solution for $V_{N}^{\beta}\left(x_{x}^{\star}(1)\right)$ based on $u_{x}^{\star}$. Therefore, we use the weak turnpike property and $\sigma(N)=\alpha^{-1}(\sqrt[-4]{N})$ to obtain (13). Let $P\in\mathbb{I}_{[0,N-1]}$ be the largest time point where $(x_{x}^{\star},u_{x}^{\star})$ is in the $\sigma(N)$-neighborhood of ${\Pi^{\star}}$, then (13) implies $P\geq N-\sqrt[4]{N^{3}}-1$. First, we consider only $N\geq N_{0}$ and choose $N_{0}\in\mathbb{N}$ large enough such that $P\geq 1$ and $\sigma(N)\leq\kappa$ for all $N\geq N_{0}$ with $\kappa$ from Assumption 7. Hence, at $P$ the point $x_{x}^{\star}(P)$ is in the local controllability neighborhood of ${\Pi^{\star}_{\mathbb{X}}}$ such that we know that there exists $u_{1}\in\mathbb{U}^{M_{1}}(x_{x}^{\star}(P))$ with $x_{u_{1}}(M_{1},x_{x}^{\star}(P))=x_{x}^{\star}(P)$. Thus, we can construct an almost optimal candidate solution $\bar{u}\in\mathbb{U}^{N}(x_{x}^{\star}(1))$ for $V_{N}^{\beta}(x_{x}^{\star}(1))$ based on the solution for $V_{N}^{\beta}(x)$ as follows
$\bar{u}(k)=\begin{cases}u_{x}^{\star}(k+1)&\text{for }k\in\mathbb{I}_{[0,\,P-2]}\\ u_{1}(k-P+1)&\text{for }k\in\mathbb{I}_{[P-1,\,P+M_{1}-2]}\\ u_{x}^{\star}(k-M_{1}+1)&\text{for }k\in\mathbb{I}_{[P+M_{1}-1,\,N-1]}.\end{cases}$ (16)
The resulting state trajectory is
$x_{\bar{u}}(k,x_{x}^{\star}(1))=\begin{cases}x_{x}^{\star}(k+1)&\text{for }k\in\mathbb{I}_{[0,P-2]}\\ x_{u_{1}}(k-P+1,x_{x}^{\star}(P))&\text{for }k\in\mathbb{I}_{[P-1,P+M_{1}-2]}\\ x_{x}^{\star}(k-M_{1}+1)&\text{for }k\in\mathbb{I}_{[P+M_{1}-1,N]}.\end{cases}$ (17)
@@
With this candidate solution we can upper bound $V_{N}^{\beta}(x_{x}^{\star}(1))\leq J_{N}^{\beta}(x_{x}^{\star}(1),\bar{u})=\Sigma_{2}+\Sigma_{3}+\Sigma_{4}$ where the three sums $\Sigma_{2}$, $\Sigma_{3}$, $\Sigma_{4}$ correspond to the three pieces in (16) and (17)
$\Sigma_{2}$ $:=\sum_{k=0}^{P-2}{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\ell\left(x_{x}^{\star}(k+1),u_{x}^{\star}(k+1)\right)$
$\Sigma_{3}$ $:=\sum_{k=P-1}^{P+M_{1}-2}{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\ell\left(x_{u_{1}}\left(k-P+1,x_{x}^{\star}(P)\right),u_{1}(k-P+1)\right)$
$\Sigma_{4}$ $:=\sum_{k=P-1}^{N-M_{1}-1}{\textstyle\beta\big{(}\frac{k+M_{1}}{N}\big{)}}\ell\left(x_{x}^{\star}(k+1),u_{x}^{\star}(k+1)\right).$
@@
$\Sigma_{3}\leq M_{1}\bar{\ell}{\textstyle\beta\big{(}\frac{P-1}{N}\big{)}}\leq M_{1}\bar{\ell}{\textstyle\beta\big{(}\frac{N-\sqrt[4]{N^{3}}-2}{N}\big{)}}=M_{1}\bar{\ell}\beta\left({\textstyle 1-\frac{1}{\sqrt[4]{N}}-\frac{2}{N}}\right)=:\delta_{7}(N)$
@@
$\Sigma_{2}+\Sigma_{4}$ $\leq\sum_{k=1}^{N-1}{\textstyle\beta\big{(}\frac{k-1}{N}\big{)}}\ell\left(x_{x}^{\star}(k),u_{x}^{\star}(k)\right)$
$=V_{N}^{\beta}(x)-\ell(x,u_{x}^{\star}(0))+\sum_{k=1}^{N-1}\left({\textstyle\beta\big{(}\frac{k-1}{N}\big{)}}-{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\right)\ell\left(x_{x}^{\star}(k),u_{x}^{\star}(k)\right)$
@@
for some $\delta_{1}\in\mathcal{L}$. Define $\delta_{2}(N):=\delta_{1}(N)+\delta_{7}(N)$ then the desired inequality (15) follows for $N\geq N_{0}$. For $N<N_{0}$, we know that (15) holds with $\delta_{2}(N):=\max\{(N_{0}+1)\bar{\ell},\delta_{2}(N_{0})\}$ and since this guarantees $\delta_{2}\in\mathcal{L}$, the desired result follows for all $N\in\mathbb{N}$.
Lemma A.7 (Difference of value functions).
Let Assumption 1-5 hold. Then, there exists $\delta_{3}\in\mathcal{L}$ such that for all $x,y\in\mathbb{X}$, $N\in\mathbb{N}$ the following inequality holds
Proof A.8.
This Lemma is the generalization of [15, Lemma 29] from linear discounts to discounts that satisfy Assumption 5. Also the proof follows the same line of arguments. First, note that if we replace the stage cost $\ell$ with the rotated stage cost $\tilde{\ell}$ from (6), then the Assumptions 15 still hold and thus Lemma A.1 applies and solutions of $\tilde{V}_{N}^{\beta}$ satisfy the weak turnpike property as well. Using the weak turnpike property (11) for $V_{N}^{\beta}(x)$, $V_{N}^{\beta}(y)$, $\tilde{V}_{N}^{\beta}(x)$, and $\tilde{V}_{N}^{\beta}(y)$ yields that there are at least $N-\frac{4\sqrt{N}}{\alpha(\varepsilon)}$ time points $k\in\mathbb{I}_{[0,N-1]}$ at which all of these four trajectories are in the $\varepsilon$ neighborhood of ${\Pi^{\star}}$. We choose $\varepsilon=\sigma_{2}(N):=\alpha^{-1}(4\sqrt[-4]{N})$ to obtain that there are at least $N-\sqrt[4]{N^{3}}$ such points. Let $P\geq N-\sqrt[4]{N^{3}}-1$ be the largest of these points and choose $N_{0}$ large enough such that $P\geq 1$ and $\sigma_{2}(N)\leq\kappa$ for $N\geq N_{0}$. First, consider only $N\geq N_{0}$, then using Assumption 7, we construct a suboptimal candidate solution $\bar{u}\in\mathbb{U}^{N}(y)$ for $\tilde{V}_{N}^{\beta}(y)$ that follows the optimal solution of $V_{N}^{\beta}(y)$ for $P$ steps, then goes in $M_{1}$ steps to the optimal orbit ${\Pi^{\star}}$ and stays there. Thus, we can bound with ~$\tilde{\ell}(\Pi^{\star}(k))=0$ (see [15, Lemma 27]) and with (7) from Assumption 7 that
$\tilde{V}_{N}^{\beta}(y)\leq\tilde{J}_{N}^{\beta}(y,\bar{u})\leq\sum_{k=0}^{P-1}{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\tilde{\ell}\left(x_{y}^{\star}(k),u_{y}^{\star}(k)\right)+\delta_{8}(N)$ (19)
with $\delta_{8}(N):=M_{1}\overline{\alpha}_{\tilde{\ell}}(\alpha_{c}(\sigma_{2}(N)))$, which is $\delta_{8}\in\mathcal{L}$. Further, since $\tilde{\ell}$ is non-negative, we can bound
$\tilde{V}_{N}^{\beta}(x)\geq\sum_{k=0}^{P-1}{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\tilde{\ell}\left(\tilde{x}_{x}^{\star}(k),\tilde{u}_{x}^{\star}(k)\right).$ (20)
@@
$\tilde{V}_{N}^{\beta}(y)-\tilde{V}_{N}^{\beta}(x)$ $\leq\delta_{8}(N)+\underbrace{\sum_{k=0}^{P-1}{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\left(\ell\left(x_{y}^{\star}(k),u_{y}^{\star}(k)\right)-\ell\left(\tilde{x}_{x}^{\star}(k),\tilde{u}_{x}^{\star}(k)\right)\right)}_{=:\Sigma_{5}}$
$\quad+\lambda(y)-\lambda(x)+\underbrace{\sum_{k=1}^{P-1}\left({\textstyle\beta\big{(}\frac{k-1}{N}\big{)}}-{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\right)\left(\lambda(\tilde{x}_{x}^{\star}(k))-\lambda(x_{y}^{\star}(k))\right)}_{=:\Sigma_{6}}$
$\quad+{\textstyle\beta\big{(}\frac{P-1}{N}\big{)}}\left(\lambda(\tilde{x}_{x}^{\star}(P))-\lambda(x_{y}^{\star}(P))\right).$ (21)
@@
${\textstyle\beta\big{(}\frac{P-1}{N}\big{)}}\left(\lambda(\tilde{x}_{x}^{\star}(P))-\lambda(x_{y}^{\star}(P))\right)\leq 2\bar{\lambda}{\textstyle\beta\big{(}\frac{P-1}{N}\big{)}}\leq 2\bar{\lambda}\beta\left({\textstyle 1-\frac{1}{\sqrt[4]{N}}-\frac{2}{N}}\right)$

### v2

@@ 5 Conclusion and outlook
$\bigg{|}\sum_{j\in\mathcal{J}}\sum_{k\in\mathcal{I}_{j}}\left({\textstyle\beta\left(\frac{k-1}{N}\right)}-{\textstyle\beta\left(\frac{k}{N}\right)}\right)g\left(x_{x}^{\star}(k),u_{x}^{\star}(k)\right)\bigg{|}\leq\left(\sqrt[4]{N^{3}}+1\right)p^{\star}\bar{g}\frac{L_{\beta}}{N}=:\delta_{4}(N)$
@@
where continuity of $g$ and compactness of $\mathbb{X}\times\mathbb{U}$ imply $\bar{g}=\max_{(x,u)\in\mathbb{X}\times\mathbb{U}}|g(x,u)|$ exists and is finite. Note that $\delta_{4}\in\mathcal{L}$. Further, with the same reasons there exists $\overline{\alpha}_{g}\in\mathcal{K}_{\infty}$ that satisfies $|g(x,u)-g({\Pi^{\star}}(k))|\leq\overline{\alpha}_{g}(\|(x,u)-{\Pi^{\star}}(k)\|)$ for all $k\in\mathbb{I}_{[0,p^{\star}-1]}$ and all $(x,u)\in\mathbb{X}\times\mathbb{U}$. Next, define the complement of $\mathcal{J}$ by $\bar{\mathcal{J}}:=\mathbb{I}_{[0,\lfloor(N-1)/p^{\star}\rfloor]}\setminus\mathcal{J}$, and compute
@@
$\Sigma_{1}$ $\leq\bigg{|}g^{\star}-\sum_{j\in\bar{\mathcal{J}}}\sum_{k\in\mathcal{I}_{j}}\left({\textstyle\beta\left(\frac{k-1}{N}\right)}-{\textstyle\beta\left(\frac{k}{N}\right)}\right)g\big{(}x_{x}^{\star}(k),u_{x}^{\star}(k)\big{)}\bigg{|}+\delta_{4}(N)$
@@
$\leq\bigg{|}g^{\star}-\sum_{j\in\bar{\mathcal{J}}}\sum_{k\in\mathcal{I}_{j}}\left({\textstyle\beta\left(\frac{k-1}{N}\right)}-{\textstyle\beta\left(\frac{k}{N}\right)}\right)g\big{(}{\Pi^{\star}}(i_{k})\big{)}\bigg{|}+\underbrace{L_{\beta}\overline{\alpha}_{g}\big{(}\sigma(N)\big{)}+\delta_{4}(N)}_{=:\delta_{5}(N)}$
@@
where $i_{k}\in\mathbb{I}_{[0,p^{\star}-1]}$ is such that $\|{\Pi^{\star}}(i_{k})-\big{(}x_{x}^{\star}(k),u_{x}^{\star}(k)\big{)}\|\leq\sigma(N)$ and $\delta_{5}\in\mathcal{L}$. First, we consider $N\geq N_{0}$ and choose $N_{0}\in\mathbb{N}$ large enough such that $\sigma(N_{0})\leq\bar{\varepsilon}$ with $\bar{\varepsilon}$ from [14, Lemma 15]. Then this Lemma guarantees that the points in $\mathcal{I}_{j}$ follow the orbit, since they are inside the $\sigma(N)$-neighborhood of ${\Pi^{\star}}$, i.e., $i_{k+1}=[i_{k}+1]_{p^{\star}}$ for all $k\in\mathcal{I}_{j}$. Hence, the $p^{\star}$ elements of $\{i_{k}|k\in\mathcal{I}_{j}\}$ are exactly $[0,p^{\star}-1]$ and for each $p\in[0,p^{\star}-1]$ and $j\in\bar{\mathcal{J}}$ there exists a $k_{p,j}\in\mathcal{I}_{j}$ such that $i_{k_{p,j}}=p$. Therefore we can continue as follows
@@
$\Sigma_{1}$ $\leq\bigg{|}\sum_{p=0}^{p^{\star}-1}g\big{(}{\Pi^{\star}}(p)\big{)}\bigg{(}\frac{1}{p^{\star}}-\sum_{j\in\bar{\mathcal{J}}}\left({\textstyle\beta\left(\frac{k_{p,j}-1}{N}\right)}-{\textstyle\beta\left(\frac{k_{p,j}}{N}\right)}\right)\bigg{)}\bigg{|}+\delta_{5}(N)$
@@
$\leq\sum_{p=0}^{p^{\star}-1}\left|g\big{(}{\Pi^{\star}}(p)\big{)}\right|\bigg{|}\frac{1}{p^{\star}}-\sum_{j\in\bar{\mathcal{J}}}\left({\textstyle\beta\left(\frac{k_{p,j}-1}{N}\right)}-{\textstyle\beta\left(\frac{k_{p,j}}{N}\right)}\right)\bigg{|}+\delta_{5}(N).$ (14)
@@
$\sum_{j\in\bar{\mathcal{J}}}\left({\textstyle\beta\left(\frac{k_{p,j}-1}{N}\right)}-{\textstyle\beta\left(\frac{k_{p,j}}{N}\right)}\right)\leq\sum_{j\in\bar{\mathcal{J}}}-\frac{1}{N}\inf_{\xi\in\left[\frac{k_{p,j}-1}{N},\frac{k_{p,j}}{N}\right]}\beta^{\prime}(\xi)$
@@
$\frac{1}{p^{\star}}-\delta_{6}(N)\leq\sum_{j\in\bar{\mathcal{J}}}\left({\textstyle\beta\left(\frac{k_{p,j}-1}{N}\right)}-{\textstyle\beta\left(\frac{k_{p,j}}{N}\right)}\right)\leq\frac{1}{p^{\star}}+\delta_{6}(N).$
@@
Using this in (14) and defining $\delta_{1}(N):=\delta_{5}(N)+\sum_{p=0}^{p^{\star}-1}\left|g({\Pi^{\star}}(p))\right|(\delta_{6}(N))$, which satisfies $\delta_{1}\in\mathcal{L}$, we obtain the desired result for $N\geq N_{0}$. For $N<N_{0}$ we know that $\Sigma_{1}\leq|g^{\star}|+\bar{g}$ and thus by redefining $\delta_{1}(N):=\max\{\delta_{1}(N_{0}),|g^{\star}|+\bar{g}\}$ for $N<N_{0}$ the result follows for all $N\in\mathbb{N}$ as still $\delta_{1}\in\mathcal{L}$. ∎
Proof A.5 (Proof of Lemma 3.3).
@@
To begin with, we construct a candidate solution for $V_{N}^{\beta}\left(x_{x}^{\star}(1)\right)$ based on $u_{x}^{\star}$. Therefore, we use the weak turnpike property and $\sigma(N)=\alpha^{-1}(\sqrt[-4]{N})$ to obtain (13). Let $P\in\mathbb{I}_{[0,N-1]}$ be the largest time point where $(x_{x}^{\star},u_{x}^{\star})$ is in the $\sigma(N)$-neighborhood of ${\Pi^{\star}}$, then (13) implies $P\geq N-\sqrt[4]{N^{3}}-1$. First, we consider only $N\geq N_{0}$ and choose $N_{0}\in\mathbb{N}$ large enough such that $P\geq 1$ and $\sigma(N)\leq\kappa$ for all $N\geq N_{0}$ with $\kappa$ from Assumption 4. Hence, at $P$ the point $x_{x}^{\star}(P)$ is in the local controllability neighborhood of ${\Pi^{\star}_{\mathbb{X}}}$ such that we know that there exists $u_{1}\in\mathbb{U}^{M_{1}}(x_{x}^{\star}(P))$ with $x_{u_{1}}(M_{1},x_{x}^{\star}(P))=x_{x}^{\star}(P)$. Thus, for $V_{N}^{\beta}(x_{x}^{\star}(1))$ we can construct a candidate solution $\bar{u}\in\mathbb{U}^{N}(x_{x}^{\star}(1))$ and the resulting state trajectory $x_{\bar{u}}$ based on the solution for $V_{N}^{\beta}(x)$ as follows
$\bar{u}(k)$ =$=\begin{cases}u_{x}^{\star}(k+1)&\text{for }k\in\mathbb{I}_{[0,\,P-2]}\\ u_{1}(k-P+1)&\text{for }k\in\mathbb{I}_{[P-1,\,P+M_{1}-2]}\\ u_{x}^{\star}(k-M_{1}+1)&\text{for }k\in\mathbb{I}_{[P+M_{1}-1,\,N-1]}.\end{cases}$ (15)
$x_{\bar{u}}(k,x_{x}^{\star}(1))$ =$=\begin{cases}x_{x}^{\star}(k+1)&\text{for }k\in\mathbb{I}_{[0,P-2]}\\ x_{u_{1}}\big{(}k-P+1,x_{x}^{\star}(P)\big{)}&\text{for }k\in\mathbb{I}_{[P-1,P+M_{1}-2]}\\ x_{x}^{\star}(k-M_{1}+1)&\text{for }k\in\mathbb{I}_{[P+M_{1}-1,N]}.\end{cases}$ (16)
@@
With this candidate solution we can upper bound $V_{N}^{\beta}\big{(}x_{x}^{\star}(1)\big{)}\leq J_{N}^{\beta}\big{(}x_{x}^{\star}(1),\bar{u}\big{)}=\Sigma_{2}+\Sigma_{3}+\Sigma_{4}$ where $\Sigma_{2}$, $\Sigma_{3}$, $\Sigma_{4}$ sum up the three pieces in (15) and (16)
$\Sigma_{2}$ $:=\sum_{k=0}^{P-2}{\textstyle\beta\left(\frac{k}{N}\right)}\ell\big{(}x_{x}^{\star}(k+1),u_{x}^{\star}(k+1)\big{)}$
$\Sigma_{3}$ $:=\sum_{k=P-1}^{P+M_{1}-2}{\textstyle\beta\left(\frac{k}{N}\right)}\ell\big{(}x_{u_{1}}\left(k-P+1,x_{x}^{\star}(P)\right),u_{1}(k-P+1)\big{)}$
$\Sigma_{4}$ $:=\sum_{k=P-1}^{N-M_{1}-1}{\textstyle\beta\left(\frac{k+M_{1}}{N}\right)}\ell\big{(}x_{x}^{\star}(k+1),u_{x}^{\star}(k+1)\big{)}.$
@@
$\Sigma_{3}\leq M_{1}\bar{\ell}{\textstyle\beta\left(\frac{P-1}{N}\right)}\leq M_{1}\bar{\ell}{\textstyle\beta\left(\frac{N-\sqrt[4]{N^{3}}-2}{N}\right)}=M_{1}\bar{\ell}\beta\left({\textstyle 1-\frac{1}{\sqrt[4]{N}}-\frac{2}{N}}\right)=:\delta_{7}(N)$
@@
$\Sigma_{2}+\Sigma_{4}$ $\leq\sum_{k=1}^{N-1}{\textstyle\beta\left(\frac{k-1}{N}\right)}\ell\left(x_{x}^{\star}(k),u_{x}^{\star}(k)\right)$
$=V_{N}^{\beta}(x)-\ell(x,u_{x}^{\star}(0))+\sum_{k=1}^{N-1}\left({\textstyle\beta\left(\frac{k-1}{N}\right)}-{\textstyle\beta\left(\frac{k}{N}\right)}\right)\ell\left(x_{x}^{\star}(k),u_{x}^{\star}(k)\right)$
@@
for some $\delta_{1}\in\mathcal{L}$. Define $\delta_{2}(N):=\delta_{1}(N)+\delta_{7}(N)$ then the desired inequality (9) follows for $N\geq N_{0}$. For $N<N_{0}$, we know that (9) holds with $\delta_{2}(N):=\max\{(N_{0}+1)\bar{\ell},\delta_{2}(N_{0})\}$ and since this guarantees $\delta_{2}\in\mathcal{L}$, the desired result follows for all $N\in\mathbb{N}$. ∎
Proof A.6 (Proof of Lemma 3.4).
This Lemma is the generalization of [17, Lemma 29] from linear discounts to discounts that satisfy Assumption 6. Also the proof follows the same line of arguments. First, note that if we replace the stage cost $\ell$ with the rotated stage cost $\tilde{\ell}$ from (5), then the Assumptions 26 still hold and thus Lemma A.1 applies and solutions of $\tilde{V}_{N}^{\beta}$ satisfy the weak turnpike property as well. The weak turnpike property (11) for $V_{N}^{\beta}(x)$, $V_{N}^{\beta}(y)$, $\tilde{V}_{N}^{\beta}(x)$, and $\tilde{V}_{N}^{\beta}(y)$ guarantees that there are at least $N-\frac{4\sqrt{N}}{\alpha(\varepsilon)}$ time points $k\in\mathbb{I}_{[0,N-1]}$ where these four trajectories are $\varepsilon$ close to ${\Pi^{\star}}$. Choosing $\varepsilon=\sigma_{2}(N):=\alpha^{-1}(4\sqrt[-4]{N})$ yields that there are at least $N-\sqrt[4]{N^{3}}$ such points. We denote the largest one by $P\geq N-\sqrt[4]{N^{3}}-1$ and choose $N_{0}$ large enough such that $P\geq 1$ and $\sigma_{2}(N)\leq\kappa$ for $N\geq N_{0}$. First, consider only $N\geq N_{0}$, then we use Assumption 4 to construct a candidate solution $\bar{u}\in\mathbb{U}^{N}(y)$ for $\tilde{V}_{N}^{\beta}(y)$ which consists of the first $P$ steps of the optimal solution of $V_{N}^{\beta}(y)$ and goes afterwards in $M_{1}$ steps to ${\Pi^{\star}}$ using the local controllability and then remains there. Using this candidate solution, we can upper bound
$\tilde{V}_{N}^{\beta}(y)\leq\tilde{J}_{N}^{\beta}(y,\bar{u})\leq\sum_{k=0}^{P-1}{\textstyle\beta\left(\frac{k}{N}\right)}\tilde{\ell}\left(x_{y}^{\star}(k),u_{y}^{\star}(k)\right)+\delta_{8}(N)$ (17)
where we used ~$\tilde{\ell}(\Pi^{\star}(k))=0$ (see [17, Lemma 27]) and (6) from Assumption 4 to compute $\delta_{8}(N):=M_{1}\overline{\alpha}_{\tilde{\ell}}(\alpha_{c}(\sigma_{2}(N)))$, which is $\delta_{8}\in\mathcal{L}$. Furthermore, non-negativity of $\tilde{\ell}$ implies
$\tilde{V}_{N}^{\beta}(x)\geq\sum_{k=0}^{P-1}{\textstyle\beta\left(\frac{k}{N}\right)}\tilde{\ell}\left(\tilde{x}_{x}^{\star}(k),\tilde{u}_{x}^{\star}(k)\right).$ (18)
@@
$\tilde{V}_{N}^{\beta}(y)-\tilde{V}_{N}^{\beta}(x)$ $\leq\delta_{8}(N)+\underbrace{\sum_{k=0}^{P-1}{\textstyle\beta\left(\frac{k}{N}\right)}\left(\ell\left(x_{y}^{\star}(k),u_{y}^{\star}(k)\right)-\ell\left(\tilde{x}_{x}^{\star}(k),\tilde{u}_{x}^{\star}(k)\right)\right)}_{=:\Sigma_{5}}$
$\quad+\lambda(y)-\lambda(x)+\underbrace{\sum_{k=1}^{P-1}\left({\textstyle\beta\left(\frac{k-1}{N}\right)}-{\textstyle\beta\left(\frac{k}{N}\right)}\right)\left(\lambda(\tilde{x}_{x}^{\star}(k))-\lambda(x_{y}^{\star}(k))\right)}_{=:\Sigma_{6}}$
$\quad+{\textstyle\beta\left(\frac{P-1}{N}\right)}\left(\lambda(\tilde{x}_{x}^{\star}(P))-\lambda(x_{y}^{\star}(P))\right).$ (19)
@@
${\textstyle\beta\left(\frac{P-1}{N}\right)}\left(\lambda(\tilde{x}_{x}^{\star}(P))-\lambda(x_{y}^{\star}(P))\right)\leq 2\bar{\lambda}{\textstyle\beta\left(\frac{P-1}{N}\right)}\leq 2\bar{\lambda}\beta\left({\textstyle 1-\frac{1}{\sqrt[4]{N}}-\frac{2}{N}}\right)$